Invariant Measures and Invariant Manifolds

Workshop on Ergodic Theory and Numerical Analysis of Dynamical Systems June 2 - 5, 1996 in Brakel, Germany

Monday 03.06.1996

09.00 Opening

09.15 H. E. Nusse, Groningen: The structure of basins of attraction and

the occurrence of Wada basins 10.15 M. Blank, Nice: Round-off pathologies and related statistics 10.45 Coffee break 11.15 P. E. Kloeden, Geelong: Minimal center of attraction of measurable

systems and their discretizations 12.15 End of morning lectures

12.30 Lunch

14.00 J. Lorenz, Albuquerque: Breakdown of invariant tori 15.00 V. Reichelt, Aachen: Invariant tori in dynamical systems 15.30 Coffee break 16.00 B. Werner, Hamburg: Numerical computation of the invariant mea-

sure for circle maps 16.30 C. Beck, London: Scaling behaviour of period lengths of dynamical

systems with roundoff 17.30 End of afternoon lectures

18.30 Dinner

20.00 M. Dellnitz, Bayreuth, M. Rumpf, Freiburg: Exploring invariant sets and invariant measures Tuesday 04.06.1996

09.00 Y. Kifer, Jerusalem: Computational aspects of random perturbations

of dynamical systems 10.00 B. Schmalfuß, Bremen: A fixed point theorem based on Lyapunov

exponents 10.30 Coffee break 11.00 A. Stuart, Stanford: Probabilistic techniques in the numerical ana-

lysis of dynamical systems 12.00 A. J. Homburg, Berlin: Eventually expanding one dimensional maps

12.30 Lunch

14.00 L. Arnold, Bremen: Invariant measures and invariant manifolds for

random dynamical systems 15.00 G. Ochs, Bremen: Stochastic stability of Oseledets spaces 15.30 Coffee break 16.00 A. Bäcker, Ulm: Symbolic dynamics and periodic orbits for the car-

dioid billiard 16.30 D. Saupe, Freiburg: The inverse problem for partitioned iterated

function systems in image compression 17.30 End of afternoon lectures

18.30 Dinner

Wednesday 05.06.1996

09.00 E. J. Doedel, Montreal: Numerical bifurcation analysis: Objectives,

methods, capability, limitations 10.00 J. Wenzelburger, Bielefeld: A statistical method for detecting cycles

in discrete dynamical systems 10.30 Coffee break 11.00 G. Keller, Erlangen: Observable invariant measures and the varia-

tional principle 12.00 End of lectures

12.15 Lunch